The ATSR Project



ATSR-2 Status :-

Under normal circumstances ATSR-2 operates under ERS-2 Mission Planning rules in a very systematic configuration. An overview of these mission planning rules, as far as they affect ATSR-2, is given in the user guide, sections 5.3 to 5.5.

If for any reason the instrument is unable to follow the mission plan then behind the following table of dates is a short description of the problem.. These events relate to the instrument status and not necessarily to data availability.


15th April 2002

ICU reset at ~07:30 UTC followed by restart of MWR and GOME and IRR cool down at ~12.40 UTC, the restart went smoothly and the instrument was commanded to NOMINAL at 17:26:40 UTC. The auto gain/offset loops operating correctly.

16th January 2002

Whilst ESOC commission EXTRA-BACKUP-MODE (an attitude control mode planned for NO-gyro operation) the instrument pointing will be degraded, these activities commenced at 016:06:30:00 UTC and are expected to be completed by 019:17:00:00 UTC.

The ATSR-2 IRR will remain on during those activities and it's expected that co-location of forward/nadir views in the data processing will be affected .

12th Apil 2002

Instrument operations request #127 prepared to request ATSR-2 instruments to STANDBY followed by an ICU reset and restart for Monday 15th April. The IRR continues running with the gains/offsets set at default values.

11th April 2002

Due to an on-board anomaly the ATSR-2 instrument suffered an emergency switch down at 101:02:18:48 UTC. Recovery of the IRR instrument began at approx. 16:00 UTC with cool down and operation complete by 23:00 UTC. However, it was noticed that the auto gain/offset loops were frozen as has been seen before.

21st March 2002

The ATSR-2 IRR detector cool down took place on 21/03/02 and the instrument operation returned to nominal at 080:18:50:41 UTC.

20th March 2002

The payload shutdown and subsequent safe mode were caused by a problem with the RA instrument ICU data-bus. The platform is now out of safe-mode. Troubleshooting the RA problem today was successful and the rest of the payload will be switched on. ATSR-2 switch on and return to operation is expected on 21/03/02.

8th March 2002

A platform anomaly occurred at 067:03:29:03 UTC which caused a payload shutdown. The reacquisition of nominal platform steering is still ongoing as of early Tuesday 12th March.

28/29th November 2001

ATSR-2 IRR reconfiguration began at 332:21:05:00 UTC. All went nominally and the instrument was nominal by 333:10:55:46 UTC.

27th November 2001

Due to ESOC planned requirements for 1 Gyro Coarse Mode Commissioning the whole payload was shut down at 331:08:39 UTC.

19th November 2001

ATSR-2 IRR reconfiguration began at 323:18:00 UTC all went smoothly and the instrument was nominal by approx. 324:00:30 UTC. The auto gain/offset loops are functioning correctly now.

During the restart the SCC cooler amplitude was increased by two command steps. This is the first SCC change since launch.

17th November 2001

ERS-2 payload shut down at 321:17:30 UTC to make safe during the Leonids meteor storm.

04th September 2001

Reconfiguration and cool down of the IRR began at 248:09:15 UTC. The instrument reconfiguration was completed by 248:16:30 UTC, however, the instrument auto gain/offset loops were seen to be 'stuck' (again).

It was agreed to leave the instrument in this state at this time after sample NRT products had been analysed to show that the default gains and offsets were producing good data with no sign of saturation..

03rd September 2001

During recovery of the RA instrument after an earlier anomaly the ERS-2 payload and ATSR-2 suffered a shut down, this occurred at approx. 247:08:45 UTC.

11th June 2001

On orbit nr. 32110 the IRR Auto gain/offset loops were disabled at 162:07:26:40 for 1023 scans whilst descending over the Arabian Peninsular. The purpose of this test was to characterise the IR channel operation over hot land surfaces.

06th June 2001

Platform AOCS using ZGM mode (Zero Gyro Mode)

01st June 2001

ESOC begin a campaign of AOCS testing to improve the YSM steering, the impact will be continued lack of H/R data for the IRR whilst global wind/wave data products are taken to be used for YSM performance characterisation.

29th May 2001

After difficutly re-installing the platform attitude control mode, ATSR-IRR finally was reconfigured on 29th May when the start up and cool down went uneventfully with a return to nominal operation at 149:15:40 UTC.

21st May 2001

Due to platform anomaly the ERS-2 payload was turned off at 141:18:12:31 UTC. Recovery will involve coarse attitude mode and 3 gyro pointing in order to re-instate EBM attitude control.

30th March 2001

ERS-2 attitude and orbit control system (AOCS) software updated with the new yaw steering control mode known as EBM-YSM. This took place around 089:10:30 UTC. Co-location of ATSR imagery should be improved.

24th March 2001

Pixel read/write anomaly at 083:05:58:16 UTC.

12th February 2001

65.7w power spike at 043:22:24:55 UTC, an isolated event.

9th February 2001

ATSR-2 was successfully restarted at 11:43 on 8 February after being off for almost three weeks. The cooldown to 80K was completed at 07:49 on 9 February and is now back to nominal operations.

The ERS-2 spacecraft attitude is operating in Extra-Backup-Mode (i.e. no gyros). This will have a significant effect on ATSR-2 images, in particular the co-location of nadir and along-track views. ESA hope to improve the performance of the attitude control as the orbit behaviour becomes more understood. The aim is to attain optimum attitude control by June in readiness for the launch of ENVISAT-1 in July. To provide data for the orbit analysis, ESA will run the AMI instrument in wind-wave mode for an extended period. This means that there will be no ATSR-2 H-Rate data until 16 February (TBC).

2nd February 2001

ESOC's efforts to establish a stable Emergency Back-up Mode (EBM) attitude control of the spacecraft platform using no gyros has not been without problems. Whilst EBM tests continue it is still not possible to operate the payload.

It is expected that return of ATSR-2 IRR operation will not be before Monday 5th February.

26th January 2001

ERS-2 payload still remains unavailable. ESOC's efforts to reconfigure the spacecraft platform have taken longer than expected.

Payload operations are not now expected to restart before the end of the weekend at the earliest.

19th January 2001

At 017:19:53:00 UTC there has been a payload shutdown.

16th January 2001

Whilst ESOC commission EXTRA-BACKUP-MODE (an attitude control mode planned for NO-gyro operation) the instrument pointing will be degraded, these activities commenced at 016:06:30:00 UTC and are expected to be completed by 019:17:00:00 UTC.

The ATSR-2 IRR will remain on during those activities and it's expected that co-location of forward/nadir views in the data processing will be affected .

5th December 2000

ERS-2 platform was in FPM between 340:13:09:46 UTC and 341:03:24:52 UTC whilst orbit control manoeuvres were executed. Co-location of fotward/nadir views in the data processing will be affected.

2nd November 2000

Platform coarse mode attitude control was reconfigured in the period 307:10:26:02 UTC through 307:17:08:26 UTC during this time YSM was replaced by FPM and consequently forward/nadir views in data processed will be affected

25th October 2000

Due to Gyro 6 becoming noisy, Fine Pointing Mode was initiated at 298:16:00:54 UTC. After platform reconfiguration Yaw Steering Mode was re-instated at 299:12:08:06 UTC.

For this period of approximately 20 hours ATSR data products will have anomalous geo-location.

12th October 2000

ERS-2 platform is now operational again using mono-gyro steering.

Reactivation of ATSR IRR began Wednesday afternoon, however, time ran out before the detector cool down could begin with an anticipated return to nominal before the last Kiruna pass for the day.

Detector cool down began at 09:00 UTC today, Thursday 12th October, and it is now expected that the instrument operation will be nominal by approximately 15.30 UTC.

7th October 2000

On Saturday 7th October at 17:30:49 UTC the ERS-2 platform suffered an attitude anomaly and, as a consequence, all the payload instruments are powered off whilst recovery takes place.

ESOC estimate ATSR should be powered up again on Wednesday 11th October, pm.

8th August 2000

Due to an ATSR software anomaly, the instrument was shut down autonomously on 08/08/2000 at 221:11:47:45 UTC. A recovery was started later in the day after an instrument power cycling and the instrument was nominal again by 222.01:27:27 UTC.

However, there will need to be a further ICU reset and instrument restart when time is available (Kiruna Ground Station is occupied by the Cluster II Launch activities on 9th August).

11th July 2000

The ATSR-IRR instrument was fully recovered to nominal operation at 17:00 UTC on 6th July 2000.

Further difficulties with the payload were encountered yesterday, 10th July 2000, at 16:32:47 UTC, when all the instruments were switched off again.

Recovery of the payload has commenced today, 11th July 2000, and it is expected that the ATSR-IRR will be back to nominal later.

30th June 2000

Since Friday 30th June at 182:06:43:00 UTC all payload instruments are switched off due to a platform anomaly.

The Problem is under investigation by ESA.

15th June 2000

Power spikes peaking 48.4w were noted during the period 167:10:00 UTC through 167:14:30 UTC on the Electronics Power Converter 1 power monitor TM F900. These spikes were associated with pixel 2002 events and a rise, peaking at approximately 14:00 UTC, of about 4 degrees in the SMU encoder temperatures.

The SMU temperatures had reverted to their nominal orbital variations by the end of the GMT day.

Scan jitter events will be noticeable in SADIST products during the period reported.

29/30th May 2000

Due to an out of plane manoeuvre, the ERS-2 spacecraft was NOT in Yaw Steering Mode (YSM), but in Fine Pointing Mode (FPM) for the period 29th May 2000 at 21:03:57 through 30th May 2000 23:53:31 UTC.

During this period, co-location of ATSR-2 forward and nadir views in SADIST products will be anomalous.

At approximately 06:50 UTC 30th May the orbit plane manoeuvre was executed, this involved the whole spacecraft performing a 90 degree yaw slew to the attitude for the manoeuvre burn and a subsequent 90 degree slew back to nominal. This manoeuvre took place during the dayside of the orbit. During the period 06:50 through 07:20 UTC a number of instrument temperature anomalies were noted. (NCR ER-NC-RAL-AT-2065 refers)

25th March 2000

The re-configuration and start-up of the ATSR-IRR was completed by 085:20:59:29 UTC. The scan mirror start-up and auto gain/offset loop operation initiated nominally.

24th March 2000

Due to an ERS-2 on board detected under-voltage in Memory Module 2 power supply, the payload was switched to STANDBY, this occurred at 084:18:15:46 UTC when there was an abrupt end to data collection.

22nd March 2000

61.2w power spike at 082:09:30:03 UTC this was an isolated case, causing a number of Pixel 2002 events, after which power consumption and SMU running was nominal.

21st March 2000

After an ICU reset and reconfiguration of the GOME and MWR instruments the IRR start-up began at approx. 081:09:40 UTC with the SMU start being nominal. The detector cool down continued at 11:20 UTC and the instrument was back to nominal operation by 18:00 UTC.

The auto gain/offset loop (recently frozen after 2 instrument restarts) kicked into operation straight away !

The instrument is now functioning in a fully nominal state.

20th March 2000

During a planned ATSR-GOME operational activity by ESOC the ATSR Instrument Control Unit (ICU) went into an unplanned status (WAIT).

ATSR payload instrument operations halted at 080:16:12:44 UTC.

9th March 2000

Checks made on low rate data from the one orbit test on the 7th March show the current settings of gain/offset in the 11 and 12 uM channels to be satisfactory for 12 bit retrieval using the 11/12uM difference coding.

12 bit difference coding was commanded at 069:12:40:52 UTC.

Trend plots over the duration of the mission have shown that 'auto' adjusted values of gain/offset drift slowly, there are also step changes in 'auto' adjusted gains and offset immediately after outgassings.

It is foreseen that the current setting will be updated from time to time and scheme of monitoring will be set up whereby these commanded values are reviewed twice monthly.

7th March 2000

Yet a further iteration at 11/12 gain/offset setting was calculated and up-linked at 067:13:41 UTC

Channel Gain Offset
11 8.76 -0.710
12 10.57 -0.262

with 12 bit difference coding for the 11/12uM channels commanded at 067:13:42 UTC

10 bit coding for the 11/12uM data was re-commanded at 067:15:18 UTC

6th March 2000

A further iteration at 12uM offset setting was calculated and up-linked at 066:12:33 UTC

Channel Gain Offset
12   -0.180

with 12 bit difference coding for the 11/12uM channels commanded at 066:12:33 UTC

10 bit coding for the 11/12uM data was re-commanded at 066:14:11 UTC

Again later analysis of the orbit of data with 12 bit coding still did not return full resolution science data.

2nd March 2000

A further iteration at gain/offset setting was calculated and up-linked at 062:16:14 UTC

Channel Gain Offset
11 8.72 -0.745
12 10.6 -0.191

with 12 bit difference coding for the 11/12uM channels commanded at 062:16:17 UTC 10 bit coding for the 11/12uM data was re-commanded at 062:17:52 UTC Again later analysis of the orbit of data with 12 bit coding still did not return full resolution science data.

1st March 2000

Further command tests after the 15th Feb. on the instrument were held off whilst ESOC continued their Mono Gyro Commissioning activities and RAL analysed in detail high science data from Kiruna.

The problem appears as an inability of the instrument on-board software to activate the auto gain/offset loop, so whilst this is further investigated attempts to set up the instrument with fixed gains and offsets together with 12bit compression for the 11/12uM channels in low rate format will be made. The following SCP settings were commanded at 061:15:08 UTC

Channel Gain Offset
11 8.56 -0.734
12 10.82 -0.191
1.6   0.0466967
.87   0.0472217
.67   0.0414476
.56   0.0251756

12 bit difference coding for the 11/12uMdata in low rate format for one orbit from 061:16:45 UTC to 061:18:22 UTC when 10 bit coding was re-commanded. Later data analysis of this orbit revealed that the 12 bit difference coding was still not returning full resolution data.

15th February 2000

With the instrument in its current status and when in low-rate format using 12bit coding for the 11/12uM channels some data is being lost.

A brief test to check the commandability of the SCP units was set up.

The signal channels were successfully commanded to fixed gains/offsets (6.25v gain;0.0v offset) for the 3.7, 11 and 12uM detectors using 12bit coding for the 11/12uM channel data and the auto gain/offset loop commanded to 'enabled', which however still remained 'frozen'. These commands took place at approx. 046:16:12 UTC and remained in operation until approx. 046:17:50 UTC when the setting were set again to:-

Channel Gain Offset
3.7 8.6 -0.115
11 8.77 -0.728
12 10.42 -0.223

with 10bit coding for 11/12uM data in the low rate format.

14th February 2000

All the ATSR instruments were commanded to standby followed by an ICU reset command at approximately 045:08:30 UTC, this was followed by an IRR cool-down at 045:13:40 UTC. The instrument was fully cooled down and back in nominal by 045:19:40 UTC.

The reset though did not have the desired effect and the auto gain/offset loop remains frozen.

11th February 2000

Data from the cool down and reconfiguration of the IRR unfortunately shows that the auto gain/offset loop for the IR channels was not operating as has occurred twice before. A request has been made to ESOC to reset the ATSR ICU, this of course requires that GOME/MWR and IRR are commanded to STANDBY and needs coordination.

10th February 2000

With early tests of the mono gyro commissioning completed and the platform in 1GYSM the ATSR-2 IRR was powered-up, configured and cool-down commenced at 041:15:30 UTC approx. The detectors reached nominal temperatures at approx. 22:00 UTC.

7th February 2000

The ERS-2 payload including the ATSR-2 IRR was shut down at the commencement of the Mono Gyro AOCS software In-Flight Commissioning. Data collection ceased at 038:06:59:00 UTC.

13th- 16th January 2000

Daily housekeeping data plots show Scan Mirror Unit temperatures F501/F508 slowly rising from 12:00 UTC on the 16th through to the end of the UTC day. A mean rise of approximately 2 deg. C was seen peaking at 02:00 UTC on the 14th, and between the times of 01:40 and 03:15 UTC some severe scan mirror jitter was noted associated with pixel 2002 events.

By the end of the day on the 16th January the elevated temperatures had settled back to nominal.

2nd January 2000

After ERS-2 platform and payload reconfiguration and the restart of the ground segment support software, collection of Low Bit Rate data recommenced at 002:15:57:02 UTC.

31st December 1999

In readiness for the year 2000 roll-over, the ERS ground segment/mission planning was to be de-activated and restarted after the year change-over.

The ATSR/IRR instrument was left running in nominal.

As part of the Y2K operations precautions the ERS-2 IDHT data collection/retrieval was paused from 365:12:23:00 UTC.

15th December 1999

66.3w power spike at 348:04:55:36 UTC this was an isolated case after which power monitoring was nominal.

16th – 18th November 1999

It was decided that as ESOC had planned to curtail payload operations during the Leonids Meteor Shower that a detector outgassing should be allowed for during the instrument shut-down.

The IRR shut-down began prior to the general Leonids instrument shut down with the detector warm-up beginning at approximately 09:00 UTC on the 16th November.

After clearing the Meteor shower trail in the early hours of the 18th Nov the IRR instrument reconfiguration and cool down began at ~14:45 UTC. The start up of the instrument was uneventful and cool down completed by 23:30 UTC with a resumption of nominal operations.

16th November 1999

IRR operations prior to the Leonids

31st May 1999

Pixel read/write anomaly at 151:19:54:25 UTC

28th May 1999

Pixel read/write anomaly at 148:15:40:17 UTC

26th May 1999

A series of high power spikes reported at:-

65.7w at 146:08:32:17 UTC

68.0w at 146:11:47:50 UTC

69.7w at 146:11:57:10 UTC

53.9w at 146:13:31:01 UTC

71.1w at 146:15:22:38 UTC

25th May 1999

66.3w power spike at 145:00:12:32 UTC again an isolated event

4th May 1999

68.5w power spike at 124:16:11:25 UTC again an isolated event.

22nd April 1999

66.3w power spike at 112:17:58:59 UTC this was an isolated case and power monitoring was nominal.

21st April 1999

Pixel read/write anomaly at 111:23:28:33UTC

4th April 1999

Several high power spike seen in the SMU power monitor these also were associated with 2002 pixel events.

70.8w at 094:09:40:32 UTC

50.5w at 094:11:29:26 UTC

65.7w at 094 12:38:17 UTC

After 13:00 UTC these jitter events subsided and power monitor reading were normal again.

31st March 1999

48.9w power spike at 090:19:30:12 UTC causing single pixel 2002 event.

20th March 1999

72.5w power spike at 079:22:06:09 UTC causing multiple pixel 2002 events for several minutes.

9th March 1999

66w power spike at 060:16:27:27 UTC. The next three days the SMU temperatures rose by approx. 1degC with a noticeable rise in mechanism power consumption.

1st March 1999

Scan Mirror Mechanism power supply experienced a non nominal power spike at 059:0650:22 UTC this was also associated with pixel 2002 events indicating severe scan jitter momentarily.

23rd November 1998

ICU reset operation executed at approx. 11:30 UTC. IRR data was not nominal between 327:11:32:42 and 327:18:12:00 UTC. The reset operation was successful and the gain/offset loops are now functioning correctly.

20th November 1998

It was again noticed that the IR channel auto gain/offset loop was ‘frozen’ as was the case on 10/11 June 1998. It was again determined that an ATSR ICU RESET was required. The sanction for this command required input from the GOME and MWR Project Team and was agreed for the 23rd Nov.

19th November 1998

Detector cool down begins at 10:40 UTC and was complete with nominal data collection by 16:56 UTC

18th November 1998

The IRR remained off due to lengthy operations to get the other ERS-2 instruments operational, the IRR restart will begin on the 19th Nov

17th November 1998

ERS-2 payload was shut down at 13:02 UTC in readiness for the Leonids Meteor Shower, the the planned sequence was actually brought forward by some hours as visual monitoring around the globe gave concern of more intense storm than originally expected.

26th October 1998

Due to an MPS error which began on 22nd October there has been NO Highrate data scheduled over land areas, return to ‘nominal’ scheduling around 23:00UTC

06th October 1998

For the first time since the 12th May 1998 when the scan mirror encoder temperature showed some significant non nominal excursions, an increase in temperature of approximately 2 deg.C was noted at about 10:45 UTC. The temperature slowly fell back to nominal values over the next 24 hours. At the same time the average power to drive the SMU rose by a noticeable amount (0.5w).

No noticeable level of scan jitter was observed and it is believed that the scan mirror bearing(s) may have created a new 'track' in the race and took the next few hours to 'run-in'.

15-16th September 1998

On the 15th Sept 1998 there was a payload anomaly associated with a command procedure for the RA instrument. A payload STANDBY command was issued by the on board computer switching all instruments to standby mode. This occurred at 258:20:58:25 UTC.

Recovery procedures for the ATSR-2 IRR commenced at 12:00 UTC on 16th Sept 1998 when visible data became available again.

The IR detector cool down was complete by 259:18:47:29 UTC at which time these data became nominal.

08th September 1998

12:30 UTC adjusted the ATSR-2 visible gains on the 0.56uM channel to 5.53, 0.67uM channel to 3.71 and the 0.87uM channel to 4.34 in Instrument Operations Request #110. These changes were made to compensate for observed gain drift over the mission so far and to optimise the control dynamic range.

15th June 1998

Shortly after the last IRR re-configuration it was noticed that the IR channel auto gain/offset loops appeared frozen at nominal fixed values.

It was decided to reset the ICU (Instrument Control Unit) software, this command though needed sanction from the GOME and MWR instrument teams.

All ATSR-2 instruments were commanded to STBY, the ICU reset issued and the IRR reconfigured. There was a period from approx. 12:30 UTC through to 18:30 UTC when data was non-nominal.

The reset was successful and immediately the gain/offset loops functioned correctly.

Data for the IR channels may need 'special consideration' from 159:16:30:00 through 166:12:30:00 UTC whilst the gains/offsets remained fixed

03rd June 1998

On the 3rd June 1998 ERS-2 suffered an anomaly associated with the attitude control Gyro #3. This caused a de-pointing of the spacecraft platform which triggered a payload EQSOL (an automatic equipment safe switch off)

This occurred at 154:14:43:53 UTC causing an abrupt end of data collection.

Subsequent operations by ESOC to recover the platform in safe attitude control with the desired yaw steering mode (YSM) took some time.

The recovery of the IRR began at about 10:00 UTC on 8th June with an ATSR-2 restart. The scan mirror start up was uneventful and data collection of visible data using pixel map 14 began more or less straight away.

The IR detector cool down began at the same time and the detectors reached nominal (cold) values at ~82K by 16:30 UTC

Did you find this information helpful? Yes No
 Give details...



Page last modified : Thursday, 06-Dec-07